OpenStudy (anonymous):

If you flip a coin 4 times, what is the probability you get at least 3 heads?

OpenStudy (anonymous):

There are 2^4 total possibilities and only 5 possibilities containing three heads e.g. {(1,1,1,0),(1,1,0,1),(1,0,1,1),(0,1,1,1),(1,1,1,1)} so 5/(2^4) = 5/16 = .3125